11 points A breathalyzer test is used by police in an area to determine whether a drivers blood alcohol conten BAC not consumed too much alcohol give a BAC reading from the breathalyzer as being above t level. Suppose that in fact, 6 6% of drivers are above the legal alcohol limit, and the police stop a driver at random sa o e h the legal limit, while 11.9% or drivers who have consumed too much legal limit. The devices not totally reliable. 5 86ofdrivers who have at alcohol will give a BAC reading below th Use this information and a theoretical group of 10,000 divers to construct a table for this testing scenario. Round all answers to four decimal places. Consumed too much alcohol to drive legally BAC reading is above legaBAC reading is below legal limit Total Yes No Total 10,000 1. What proportion of drivers will have a BAC reading showing them to be above the legal limit? 2. Suppose that a randomly stopped driver has a reading above the legal limit. What is the probability that they have consumed too much alcohol to drive safely? 3. What percentage of drivers who test below the legal limit will actually have consumed too much alcohol to drive safely?
Answer #1
above limit below limit total
Yes 581.4600 78.5400 660
No 541.7200 8798.2800 9340
total 1123.1800 8876.8200 10000

1)P(BAC reading above legal limit)=1123.18/10000=0.1123

2)P(too much alcohol| reading above limit) =581.46/1123.18=0.5177

3)

percentage (too much alcohol| reading below limit)=(78.54/8876.82)*100=0.8848%